OpenStudy (anonymous):

Let y = 3t + 6 be a linear function representing the distance from home for an ant t minutes after starting out from a location near its home. What does the number 3 represent in this function? A. The ant is 3 feet from its home after t minutes. B. The ant started out 3 feet from its home. C. The ant is crawling at 3 feet per minute. D. The ant is moving 3 feet every 6 minutes. ***Please explain!! :) THanks :)

jimthompson5910 (jim_thompson5910):

A table might help here

jimthompson5910 (jim_thompson5910):

t | y ---------------- 0 | 6 1 | 9 2 | 12 3 | 15 4 | 18 5 | 21 Notice how when t = 2 for instance, y is y = 3t+6 y = 3(2)+6 y = 12 So in the row that starts with 2, you write '12' in the 'y' column.

jimthompson5910 (jim_thompson5910):

So in the beginning (t = 0 minutes), the ant is 6 ft from home. 1 minute later at t = 1, the ant is 9 ft from home. The next minute after that (t = 2), the ant is 12 ft from home. etc etc

OpenStudy (anonymous):

okay i see :) so they're asking what the number 3 represents... in 3t right?

jimthompson5910 (jim_thompson5910):

yes, using the info above, what does it mean?

OpenStudy (anonymous):

the ant started 3 feet from its home?? because 3(3)+6=9+6=15??

jimthompson5910 (jim_thompson5910):

no, it started 6 ft from his home because when t = 0, y = 6

jimthompson5910 (jim_thompson5910):

its in the first row of the table

OpenStudy (anonymous):

is the ant moving 3 feet for every 6 minutes then? idk :/

OpenStudy (anonymous):

C. The ant is crawling at 3 feet per minute. because as jim says 6 is the starting distance from home.........each value of t is multiplied by 3....

jimthompson5910 (jim_thompson5910):

no, it moves 3 ft in one minute

jimthompson5910 (jim_thompson5910):

so its speed is 3 ft/min

OpenStudy (anonymous):

ohhh so the answer is C. the ant is crawling at 3 feet per minute? =) cuz y is the feet moved... and x represents the minutes??

jimthompson5910 (jim_thompson5910):

exactly

OpenStudy (anonymous):

yay :) thanks!! :)

jimthompson5910 (jim_thompson5910):

yw